Solve the following equation for x.
185.45=32.1e^1.23x

Answer by ankor@dixie-net.com(22740)

185.45 = 32.1 * e^1.23x
Or
e^1.23x = 185.45/32.1; divided both sides by 32.1
e^1.23x = 5.77725856
:
Find the nat log of both sides,
ln(e^1.23x) = ln(5.77725856)
:
The log equiv of exponents:
1.23x * ln(e) = ln(577725856)
1.23x = 1.753929; remember the ln of e is 1
x = 1.753929/1.23
x = 1.42596 (rounded)
:
Check our solution on a good calc like a Ti83: enter 32.1* (e^(1.23*1.42596)
Should get 185.45
:
Could you follow this OK? Any questions?